# Nonharmonic multivariate Fourier transforms and matrices: condition numbers and hyperplane geometry

### v1

@@ 3 Super-resolution regime and arbitrary nodes
Now, we discuss the main differences between Theorems 3.9 and 3.10. The primary difference is the metric. When Fourier samples are collected in a ball, all distances are measured in $\ell^{2}$, which is natural since the ball $B_{m}$ and the $\ell^{2}$ distance are rotationally symmetric. However, when Fourier samples are collected in a cube, while the local sparsity $\nu$ is defined using the $\ell^{\infty}$ distance, all small scales are measured in $\ell^{1}$. This is natural since $\ell^{1}$ and $\ell^{\infty}$ are dual. It also shows that taking measurements in a cube offers an advantage over the ball because $|x|_{2}\leq|x|_{1}$, so elements in $\mathcal{X}$ are more separated in $\ell^{1}$ as opposed to $\ell^{2}$. There are some slight differences in the requirements for $m$ and $\tau$, which are equivalent up to constants that depend on $p$ and $d$.
For the ball case, the constants that appear before the product terms are more complicated. Results for the generalized Gauss circle problem tell us that, for fixed dimension $d\geq 2$, the leading order growth rates of $|B_{r}|$ and $|B_{r}|_{*}$ are identical in the limit $r\to\infty$. See [28] for a survey on this topic. Consequently, $|B_{\alpha/\tau}|/|B_{\alpha/\tau}|_{*}$ can be approximated by 1 when $\alpha/\tau$ is large. There are numerous open and difficult questions about how tight this approximation is. In several examples that we discuss later on, $\alpha/\tau$ can be selected proportional to $m$, so this quotient term can be effectively ignored when $m$ is large. Such a ratio does not appear in the cube case because $|Q_{r+1}|=|Q_{r}|_{*}$ whenever $r\in\mathbb{N}$. Going back to the ball case, the constant $c(\alpha)$ is small when $d$ is large. This can be addressed using an alternative result (that is not included in this paper) for the well separated case for samples in a ball, see the discussion following Theorem 2.5.
@@
$\sigma_{\min}(\mathcal{F}_{B_{m},\mathcal{X}})$ $\geq\frac{1}{\sqrt{s}}\Big{(}\frac{|B_{1}|}{c(\alpha)}\Big{)}^{\lambda/2}\Big{(}\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}\Big{)}^{\lambda/2}\,\sqrt{|B_{m/(2\lambda)}|}\,\,\,\Big{(}\frac{m\delta}{\lambda}\Big{)}^{\lambda-1},$
$\sigma_{\min}(\Phi_{B_{m},\mathcal{X}})$ $\geq\frac{1}{\sqrt{s}}\Big{(}\frac{|B_{1}|}{c(\alpha)}\Big{)}^{\lambda/2}\Big{(}\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}\Big{)}^{\lambda/2}\sqrt{|B_{m/(2\lambda)}|_{*}}\,\,\Big{(}\frac{m\delta}{\lambda}\Big{)}^{\lambda-1}.$
@@
$\sigma_{\min}(\mathcal{F}_{Q_{m},\mathcal{X}})$ $\geq\sqrt{\frac{2}{s}}\,\Big{(}1-\frac{1}{2}e^{1/(2\beta)}\Big{)}^{\lambda/2}\,\sqrt{|Q_{m/(2\lambda)}|}\,\,\,\Big{(}\frac{m\delta}{\lambda}\Big{)}^{\lambda-1},$
$\sigma_{\min}(\Phi_{Q_{m},\mathcal{X}})$ $\geq\sqrt{\frac{2}{s}}\,\Big{(}1-\frac{1}{2}e^{1/(2\beta)}\Big{)}^{\lambda/2}\,\sqrt{|Q_{m/(2\lambda)}|_{*}}\,\,\Big{(}\frac{m\delta}{\lambda}\Big{)}^{\lambda-1}.$
The above dependence on $m$ and $\delta$ are optimal. For example, suppose the elements of $\mathcal{X}$ are equally spaced by $\delta$ on a line parallel to one of the coordinate axes. Then the matrix $\Phi_{Q_{m},\mathcal{X}}$ consists of $(2m+1)^{d-1}$ copies of a one-dimensional Fourier matrix corresponding to nodes $\{0,\delta,\dots,(\lambda-1)\delta\}$ and frequencies $\{-m,\dots,m\}$. Upper bounds for the smallest singular value of one-dimensional Fourier matrices, see [35, Proposition 2.10] tell us that the smallest singular value of each one is upper bounded by $C_{\lambda}\sqrt{m}(m\delta)^{\lambda-1}$. Using this for each of the $(2m+1)^{d-1}$ repeated matrices yields the upper bound
@@
A similar result to Corollary 3.12 was established in Kunis-Nagal [30]. They lower bounded $\sigma_{\min}(\Phi_{Q_{m}^{\prime},\mathcal{X}})$ where $Q_{m}^{\prime}=\{0,\dots,m\}^{d}$ and obtained the same $C_{\lambda,d}m^{d/2}(m\delta)^{\lambda-1}$ lower bound up (with different implicit constant), but under significantly stronger assumptions on the clump separation $\tau$. Indeed, their weakest condition on $\tau$, see [30, Examples 4.3 (ii)], requires
@@
As the minimum separation goes to zero, necessarily $\mathfrak{C}\to\infty$, which implies $\tau\to\infty$. So this condition becomes vacuous if any two nodes in $\mathcal{X}$ are pushed closer together, while the rest remain fixed. This is a proof artifact that can be traced back to the interpolation techniques developed in [35]. To compare, the separation condition in Corollary 3.12 does not depend on $\delta$, scales linearly with $\lambda$ instead of quadratically, and we obtain the same dependence on $d$. Another main difference is that they use the $\ell^{\infty}$ distance on $\mathcal{X}$, whereas our corollary uses the $\ell^{1}$ metric, which is preferable since $|x|_{\infty}\leq|x|_{1}$. Our improved results for multidimensional clumps are largely due to new technical developments for the one-dimensional problem provided by [34], which dramatically improved upon [35].
@@ 4 Super-resolution regime and hyperplane geometry
$\sigma_{\min}(\mathcal{F}_{B_{m},\mathcal{X}})$ $\geq\frac{1}{\sqrt{s}}\Big{(}\frac{|B_{1}|}{c(\alpha)}\Big{)}^{\nu/2}\Big{(}\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}\Big{)}^{\nu/2}\,\sqrt{|B_{m/(2r+2)}|}\,\,\frac{1}{2^{r/2}}\Big{(}\frac{2m\eta}{r+1}\Big{)}^{r}.$
@@
$\sigma_{\min}(\mathcal{F}_{Q_{m},\mathcal{X}})$ $\geq\sqrt{\frac{2}{s}}\,\Big{(}1-\frac{1}{2}e^{1/(2\beta)}\Big{)}^{\nu/2}\,\sqrt{|Q_{m/(2r+2)}|}\,\,\frac{1}{2^{r/2}}\Big{(}\frac{2m\eta}{r+1}\Big{)}^{r}.$
@@
Here we make some basic comparisons between Theorems 4.2 and 3.9 for the ball, and Theorems 4.3 and 3.10 for the cube. These theorem share similar assumptions and constants that depend on $\tau$, $\nu$, and $\alpha$ or $\beta$, which come from reducing the global problem into to a local one via Lemma 3.6. The main differences between these two pairs of theorems is how the local problem is handled.
@@
    Vertices of a triangle. Say $\mathcal{X}=\{(0,0),(\delta,0),(0,\delta)\}$ for sufficiently small $\delta$. Here, we see that the theorems in Section 3 predict a rate of $\delta^{2}$. On the other hand, $\mathcal{X}$ locally consists of hyperplanes with parameters $(C\delta,p,1,\delta/2)$. Indeed, for any $x\in\mathcal{X}$, there is a single line that contains $\mathcal{X}\setminus\{x\}$ and its distance to $x$ is at least $\delta/\sqrt{2}$, see Figure 4. Thus, the theorems in this section give the correct $\delta$ rate. This type of reasoning extends to any arbitrary triangle whose side lengths are on the order of $\delta$. This example was also studied in [30, Example 5.3] and the $\delta$ rate for the smallest singular value was derived via explicit calculation, whereas we obtained this using hyperplane methods that can deal with more general sets.
@@
Let us discuss some related work. Of course, it is well known that in higher dimensions, the geometry of $\mathcal{X}$ greatly influences the behavior of polynomial interpolation. Related hyperplane decompositions have been exploited in the context of Lagrange interpolation and appears in books, see [15, Chapter 10] and [41, Chapter 5]. Many prior works focus on finding polynomials of the smallest degree for which interpolation of arbitrary values on $\mathcal{X}$ is possible. Our focus is in the opposite direction. We allow for large $m$ relative to the local sparsity of $\mathcal{X}$, which corresponds to using a larger class of functions than necessary for interpolation. In this setting, there is a localization effect of the Fourier transform, so only the local (as opposed to global), hyperplane geometry of $\mathcal{X}$ predominantly affects $\sigma_{\min}(\mathcal{F}_{\Omega,\mathcal{X}})$.
@@ 5 Proofs of theorems
$\widehat{\psi_{m}}(\xi)=\psi(m^{-1}\xi)\quad\text{and}\quad\widehat{\varphi_{m}}(\xi)=\varphi(m^{-1}\xi).$
@@
$\varphi_{m}(0)$ $=r^{d}\varphi(0)\leq r^{d}e^{1/(2\beta)}|Q_{1}|=e^{1/(2\beta)}|Q_{m}|.$
@@
Now we focus on $\Phi_{Q_{m},\mathcal{X}}$. We assume without loss of generality that $m\in\mathbb{N}_{+}$ since the conclusion remains unchanged if $m$ is not an integer in view of the fact that $|Q_{m}|_{*}=|Q_{\lfloor m\rfloor}|_{*}$ and $\Phi_{Q_{m},\mathcal{X}}=\Phi_{Q_{\lfloor m\rfloor},\mathcal{X}}$. For convenience, set $r=m+1-\varepsilon$ where $\varepsilon\in(0,1)$ will be made small later. Define the dilated functions $\psi_{r}$ and $\varphi_{r}$. Since $r\geq m$, we see that $\varphi_{r}$ and $\psi_{r}$ are supported in $Q_{\beta d/r}\subseteq Q_{\beta d/m}$, that $\Delta_{\infty}(\mathcal{X})\geq\frac{\beta d}{m}\geq\frac{\beta d}{r}$, and that
@@
Combining these observations, making $\varepsilon$ arbitrarily small, and noting that $|Q_{m}|_{*}=|Q_{m+1}|=2^{d}(m+1)^{d}$ completes the proof.
@@
$\sigma_{\min}(\Phi_{B_{\alpha/\tau},\mathcal{V}})\geq\sqrt{c(\alpha)\Big{(}\frac{\alpha}{\tau}\Big{)}^{d}}=\sqrt{\frac{c(\alpha)}{|B_{1}|}\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}}\sqrt{|B_{\alpha/\tau}|_{*}}.$
@@
$\|g_{k}\|_{L^{\infty}(\mathbb{T}^{d})}\leq\Big{(}\frac{|B_{1}|}{c(\alpha)}\Big{)}^{\nu/2}\Big{(}\frac{|B_{\alpha/\tau}|_{*}}{|B_{\alpha/\tau}|}\Big{)}^{\nu/2}.$
For $b_{k}$, we will use Lemma 3.8 where the neighborhood set $\mathcal{U}=\mathcal{N}_{\infty}(x_{k},\tau,\mathcal{X})-x_{k}$, $n=\frac{m}{2}$, $r=\nu$, and $p=2$. By the theorem’s assumption that $\tau\leq\frac{1}{4\sqrt{d}}$, for any $x_{j}\in\mathcal{N}_{2}(x_{k},\tau,\mathcal{X})$, we have
@@
$h(x)=\frac{1}{|B_{m/(2\nu)}|_{*}}\sum_{\omega\in B_{m/(2\nu)}}e^{2\pi i\omega\cdot x}\quad\text{and}\quad\|h\|_{L^{2}(\mathbb{T}^{d})}=\frac{1}{\sqrt{|B_{m/(2\nu)}|_{*}}}.$
@@
$\|f_{k}\|_{L^{2}}\leq\|h\|_{L^{2}}\|b_{k}\|_{L^{\infty}}\|g_{k}\|_{L^{\infty}}\leq\|h\|_{L^{2}}\Big{(}\frac{c(\alpha)}{|B_{1}|}\Big{)}^{\nu/2}\Big{(}\frac{|B_{\alpha/\tau}|_{*}}{|B_{\alpha/\tau}|}\Big{)}^{\nu/2}\prod_{0<|x_{j}-x_{k}|_{2}\leq\frac{\nu}{m}}\,\frac{\nu}{m|x_{j}-x_{k}|_{2}}.$
@@
Both theorems have the same structure, so we prove them simultaneously. Let $p\in\{2,\infty\}$. The starting point of this proof is similar to that of Theorem 3.9 and Theorem 3.10. For each $k\in\{1,\dots,s\}$, we decompose $\mathcal{X}$ into a neighborhood set $\mathcal{N}_{p}(x_{k},\tau,\mathcal{X})$ and its complement. Let $g_{k}\in\mathcal{P}(\Omega^{p}_{m/2})$ be the function constructed in Section 5.3 for $p=2$ and Section 5.3 for $p=\infty$. Recall the previously obtained estimates for $\|g_{k}\|_{L^{\infty}(\mathbb{T}^{d})}$. Now we extending $g_{k}$ periodically to an element of $L^{\infty}(\mathbb{R}^{d})$
The neighborhood set $\mathcal{N}_{p}(x_{k},\tau,\mathcal{X})$, by assumption consists of at most $r$ hyperplanes relative to $x_{k}$ whose distances to $x_{k}$ are at least $\eta$. Applying Lemma 4.1 with $n=\frac{m}{2}$, which is justified by the assumption that $\eta\leq\frac{r+1}{2m}$, we obtain a $b_{k}\in\mathcal{W}(\Omega^{p}_{m/2})$ such that $b_{k}(x_{k})=1$, $b_{k}$ vanishes on $\mathcal{N}_{p}(x_{k},\tau,\mathcal{X})\setminus\{x_{k}\}$ and
$\|b_{k}\|_{L^{2}(\mathbb{R}^{d})}\leq\sqrt{\frac{2^{r}}{|\Omega^{p}_{m/(2r+2)}|}}\,\Big{(}\frac{r+1}{2m\eta}\Big{)}^{r}.$

### v2

@@ 3 Super-resolution regime and arbitrary nodes
Now, we discuss the main differences between Theorems 3.9 and 3.10. The primary difference is the metric. When Fourier samples are collected in a ball, all distances are measured in $\ell^{2}$, which is natural since the ball $B_{m}$ and the $\ell^{2}$ distance are rotationally invariant. However, when Fourier samples are collected in a cube, the local sparsity $\nu$ is defined using the $\ell^{\infty}$ distance, while all small scales are measured in $\ell^{1}$. This is natural since $\ell^{1}$ and $\ell^{\infty}$ are dual. It also shows that taking measurements in a cube offers an advantage over the ball because $|x|_{2}\leq|x|_{1}$, so elements in $\mathcal{X}$ are more separated in $\ell^{1}$ as opposed to $\ell^{2}$. There are some slight differences in the requirements for $m$ and $\tau$, which are equivalent up to constants that depend on $p$ and $d$.
For the ball case, the constants that appear before the product terms are more complicated. Results for the generalized Gauss circle problem tell us that, for fixed dimension $d\geq 2$, the leading order growth rates of $|B_{r}|$ and $|B_{r}|_{*}$ are identical in the limit $r\to\infty$. See [28] for a survey on this topic. Consequently, $|B_{\alpha/\tau}|/|B_{\alpha/\tau}|_{*}$ can be approximated by 1 when $\alpha/\tau$ is large. There are numerous open and difficult questions about how tight this approximation is. In several examples that we discuss later on, $\alpha/\tau$ can be selected proportional to $m$, so this quotient term can be effectively ignored when $m$ is large. The constant $c(\alpha)$ is small when $d$ is large. This can be addressed using an alternative result (that is not included in this paper) for the well separated case for samples in a ball, see the discussion following Theorem 2.5.
@@
$\sigma_{\min}(\mathcal{F}_{B_{m},\mathcal{X}})$ $\geq\frac{1}{\sqrt{2^{\lambda-1}s}}\left(\frac{c(\alpha)}{|B_{1}|}\right)^{\lambda/2}\left(\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}\right)^{\lambda/2}\,\sqrt{|B_{m/(2\lambda)}|}\,\,\,\left(\frac{m\delta}{\lambda}\right)^{\lambda-1},$
$\sigma_{\min}(\Phi_{B_{m},\mathcal{X}})$ $\geq\frac{1}{\sqrt{2^{\lambda-1}s}}\left(\frac{c(\alpha)}{|B_{1}|}\right)^{\lambda/2}\left(\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}\right)^{\lambda/2}\sqrt{|B_{m/(2\lambda)}|_{*}}\,\,\left(\frac{m\delta}{\lambda}\right)^{\lambda-1}.$
@@
$\sigma_{\min}(\mathcal{F}_{Q_{m},\mathcal{X}})$ $\geq\sqrt{\frac{2}{s}}\,\left(1-\frac{1}{2}e^{1/(2\beta)}\right)^{\lambda/2}\,\sqrt{|Q_{m/(2\lambda)}|}\,\,\,\left(\frac{m\delta}{\lambda}\right)^{\lambda-1},$
$\sigma_{\min}(\Phi_{Q_{m},\mathcal{X}})$ $\geq\sqrt{\frac{2}{s}}\,\left(1-\frac{1}{2}e^{1/(2\beta)}\right)^{\lambda/2}\,\sqrt{|Q_{m/(2\lambda)}|_{*}}\,\,\left(\frac{m\delta}{\lambda}\right)^{\lambda-1}.$
The above dependence on $m$ and $\delta$ are optimal. For example, suppose the elements of $\mathcal{X}$ are equally spaced by $\delta$ on a line parallel to one of the coordinate axes. Then the matrix $\Phi_{Q_{m},\mathcal{X}}$ consists of $(2m+1)^{d-1}$ copies of a one-dimensional Fourier matrix corresponding to nodes $\{0,\delta,\dots,(\lambda-1)\delta\}$ and frequencies $\{-m,\dots,m\}$. Upper bounds for the smallest singular value of one-dimensional Fourier matrices, see [36, Proposition 2.10] tell us that the smallest singular value of each one is upper bounded by $C_{\lambda}\sqrt{m}(m\delta)^{\lambda-1}$. Using this for each of the $(2m+1)^{d-1}$ repeated matrices yields the upper bound
@@
A similar result to Corollary 3.12 was established in Kunis-Nagal [31]. They lower bounded $\sigma_{\min}(\Phi_{Q_{m}^{\prime},\mathcal{X}})$ where $Q_{m}^{\prime}=\{0,\dots,m\}^{d}$ and obtained the same $C_{\lambda,d}m^{d/2}(m\delta)^{\lambda-1}$ lower bound up (with different implicit constant), but under significantly stronger assumptions on the clump separation $\tau$. Indeed, their weakest condition on $\tau$, see [31, Examples 4.3 (ii)], requires
@@
As the minimum separation goes to zero, necessarily $\mathfrak{C}\to\infty$, which implies $\tau\to\infty$. So this condition becomes vacuous if any two nodes in $\mathcal{X}$ are pushed closer together, while the rest remain fixed. This is a proof artifact that can be traced back to the interpolation techniques developed in [36]. To compare, the separation condition in Corollary 3.12 does not depend on $\delta$, scales linearly with $\lambda$ instead of quadratically, and has the same dependence on $d$. Another main difference is that they use the $\ell^{\infty}$ distance on $\mathcal{X}$, whereas our corollary uses the $\ell^{1}$ metric, which is preferable since $|x|_{\infty}\leq|x|_{1}$. Our improved results for multidimensional clumps are largely due to new technical developments for the one-dimensional problem provided by [35], which dramatically improved upon [36].
@@ 4 Super-resolution regime and hyperplane geometry
$\sigma_{\min}(\mathcal{F}_{B_{m},\mathcal{X}})$ $\geq\frac{1}{\sqrt{2^{\nu-1}s}}\left(\frac{c(\alpha)}{|B_{1}|}\right)^{\nu/2}\left(\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}\right)^{\nu/2}\,\sqrt{|B_{m/(2r+2)}|}\,\,\frac{1}{2^{r/2}}\left(\frac{2m\eta}{r+1}\right)^{r}.$
@@
$\sigma_{\min}(\mathcal{F}_{Q_{m},\mathcal{X}})$ $\geq\sqrt{\frac{2}{s}}\,\left(1-\frac{1}{2}e^{1/(2\beta)}\right)^{\nu/2}\,\sqrt{|Q_{m/(2r+2)}|}\,\,\frac{1}{2^{r/2}}\left(\frac{2m\eta}{r+1}\right)^{r}.$
@@
Here we make some basic comparisons between Theorems 4.2 and 3.9 for the ball, and Theorems 4.3 and 3.10 for the cube. These theorems share similar assumptions and constants that depend on $\tau$, $\nu$, and $\alpha$ or $\beta$, which come from reducing the global problem into to a local one via Lemma 3.6. The main difference between these two pairs of theorems is how the local problem is handled.
@@
    Vertices of a triangle. Say $\mathcal{X}=\{(0,0),(\delta,0),(0,\delta)\}$ for sufficiently small $\delta$. Here, we see that the theorems in Section 3 predict a rate of $\delta^{2}$. On the other hand, $\mathcal{X}$ locally consists of hyperplanes with parameters $(C\delta,p,1,\delta/2)$. Indeed, for any $x\in\mathcal{X}$, there is a single line that contains $\mathcal{X}\setminus\{x\}$ and its distance to $x$ is at least $\delta/\sqrt{2}$, see Figure 4. Thus, the theorems in this section give the correct $\delta$ rate. This type of reasoning extends to any arbitrary triangle whose side lengths are on the order of $\delta$. This example was also studied in [31, Example 5.3] and the $\delta$ rate for the smallest singular value was derived via explicit calculation, whereas we obtained this using hyperplane methods that can deal with more general sets.
@@
Let us discuss some related work. Of course, it is well known that in higher dimensions, the geometry of $\mathcal{X}$ greatly influences the behavior of polynomial interpolation. Related hyperplane decompositions have been exploited in the context of Lagrange interpolation and appears in books, see [15, Chapter 10] and [43, Chapter 5]. Many prior works focus on finding polynomials of the smallest degree for which interpolation of arbitrary values on $\mathcal{X}$ is possible. Our focus is in the opposite direction. We allow for large $m$ relative to the local sparsity of $\mathcal{X}$, which corresponds to using a larger class of functions than necessary for interpolation. In this setting, there is a localization effect of the Fourier transform, so only the local (as opposed to global), hyperplane geometry of $\mathcal{X}$ predominantly affects $\sigma_{\min}(\mathcal{F}_{\Omega,\mathcal{X}})$.
@@ 5 Proofs of theorems
$\widehat{\psi_{m}}(\xi)=\widehat{\psi}(m^{-1}\xi)\quad\text{and}\quad\widehat{\varphi_{m}}(\xi)=\widehat{\varphi}(m^{-1}\xi).$
@@
$\varphi_{m}(0)$ $=m^{d}\varphi(0)\leq m^{d}e^{1/(2\beta)}|Q_{1}|=e^{1/(2\beta)}|Q_{m}|.$
@@
Now we focus on $\Phi_{Q_{m},\mathcal{X}}$. We assume without loss of generality that $m\in\mathbb{N}_{+}$ since the conclusion remains unchanged if $m$ is not an integer in view of the fact that $|Q_{m}|_{*}=|Q_{\lfloor m\rfloor}|_{*}$ and $\Phi_{Q_{m},\mathcal{X}}=\Phi_{Q_{\lfloor m\rfloor},\mathcal{X}}$. For convenience, set $r=m+1-\varepsilon$ where $\varepsilon\in(0,1)$ will be adjusted later. Define the dilated functions $\psi_{r}$ and $\varphi_{r}$. Since $r\geq m$, we see that $\varphi_{r}$ and $\psi_{r}$ are supported in $Q_{\beta d/r}\subseteq Q_{\beta d/m}$, that $\Delta_{\infty}(\mathcal{X})\geq\frac{\beta d}{m}\geq\frac{\beta d}{r}$, and that
@@
Picking $\varepsilon=1/2$ and using that $|Q_{m}|_{*}=(2m+1)^{d}$, we obtain
$\sigma_{\min}(\Phi_{Q_{m},\mathcal{X}})$ $\geq\sqrt{\psi_{r}(0)}\geq\sqrt{(2-e^{1/(2\beta)})2^{d}(m+1/2)^{d}}=\sqrt{(2-e^{1/(2\beta)})|Q_{m}|_{*}},$
$\sigma_{\max}(\Phi_{Q_{m},\mathcal{X}})$ $\leq\sqrt{\varphi_{r}(0)}\leq\sqrt{e^{1/(2\beta)}2^{d}(m+1/2)^{d}}=\sqrt{e^{1/(2\beta)}|Q_{m}|_{*}}.$
@@
$\sigma_{\min}(\Phi_{B_{\alpha/\tau},\mathcal{V}})\geq\sqrt{c(\alpha)\left(\frac{\alpha}{\tau}\right)^{d}}=\sqrt{\frac{c(\alpha)}{|B_{1}|}\frac{|B_{\alpha/\tau}|}{|B_{\alpha/\tau}|_{*}}}\sqrt{|B_{\alpha/\tau}|_{*}}.$
@@
$\|g_{k}\|_{L^{\infty}(\mathbb{T}^{d})}\leq\left(\frac{|B_{1}|}{c(\alpha)}\right)^{\nu/2}\left(\frac{|B_{\alpha/\tau}|_{*}}{|B_{\alpha/\tau}|}\right)^{\nu/2}.$
For $b_{k}$, we will use Lemma 3.8 where the neighborhood set $\mathcal{U}=\mathcal{N}_{2}(x_{k},\tau,\mathcal{X})-x_{k}$, $n=\frac{m}{2}$, $r=\nu$, and $p=2$. By the theorem’s assumption that $\tau\leq\frac{1}{4\sqrt{d}}$, for any $x_{j}\in\mathcal{N}_{2}(x_{k},\tau,\mathcal{X})$, we have
@@
$h(x)=\frac{1}{|B_{m/(2\nu)}|_{*}}\sum_{\omega\in B_{m/(2\nu)}\cap\mathbb{Z}^{d}}e^{2\pi i\omega\cdot x}\quad\text{and}\quad\|h\|_{L^{2}(\mathbb{T}^{d})}=\frac{1}{\sqrt{|B_{m/(2\nu)}|_{*}}}.$
@@
$\|f_{k}\|_{L^{2}}$ $\leq\|h\|_{L^{2}}\|b_{k}\|_{L^{\infty}}\|g_{k}\|_{L^{\infty}}$
$\leq\|h\|_{L^{2}}\sqrt{2^{\nu-1}}\left(\frac{|B_{1}|}{c(\alpha)}\right)^{\nu/2}\left(\frac{|B_{\alpha/\tau}|_{*}}{|B_{\alpha/\tau}|}\right)^{\nu/2}\prod_{0<|x_{j}-x_{k}|_{2}\leq\frac{\nu}{m}}\,\frac{\nu}{m|x_{j}-x_{k}|_{2}}.$
@@
Both theorems have the same structure, so we prove them simultaneously. Let $p\in\{2,\infty\}$. The starting point of this proof is similar to that of Theorem 3.9 and Theorem 3.10. For each $k\in\{1,\dots,s\}$, we decompose $\mathcal{X}$ into a neighborhood set $\mathcal{N}_{p}(x_{k},\tau,\mathcal{X})$ and its complement. Let $g_{k}\in\mathcal{P}(\Omega^{p}_{m/2})$ be the function constructed in Section 5.3 for $p=2$ and Section 5.3 for $p=\infty$. Recall the previously obtained estimates for $\|g_{k}\|_{L^{\infty}(\mathbb{T}^{d})}$. Now we extend $g_{k}$ periodically to an element of $L^{\infty}(\mathbb{R}^{d})$.
By assumption, the neighborhood set $\mathcal{N}_{p}(x_{k},\tau,\mathcal{X})$ consists of at most $r$ hyperplanes relative to $x_{k}$ whose distances to $x_{k}$ are at least $\eta$. Applying Lemma 4.1 with $n=\frac{m}{2}$, which is justified by the assumption that $\eta\leq\frac{r+1}{2m}$, we obtain a $b_{k}\in\mathcal{W}(\Omega^{p}_{m/2})$ such that $b_{k}(x_{k})=1$, $b_{k}$ vanishes on $\mathcal{N}_{p}(x_{k},\tau,\mathcal{X})\setminus\{x_{k}\}$ and
$\|b_{k}\|_{L^{2}(\mathbb{R}^{d})}\leq\sqrt{\frac{2^{r}}{|\Omega^{p}_{m/(2r+2)}|}}\,\left(\frac{r+1}{2m\eta}\right)^{r}.$